OpenStudy (anonymous):

solve using elimination method: 3x+11y=34.5 9x-7y=-16.5

OpenStudy (anonymous):

3x + 11y = 34.5...................1 9x – 7y = -16.5...................2 .. multiply (1) by -3 -9x-33y=-103.5 add this to (2) -7y-33=-16.5-103.5 -40y=-120 /-40 y =3 .. plug value of y in 1 3x+33=34.5 3x=34.5-33 3x=1.5 x=1.5/3 x=0.5

OpenStudy (anonymous):

times the top by 3 ----> 9x+33y=103.5 - 9x-7y=-16.5 ---------------- 0x +40y=120 40y=120 y=120/40 y=3

OpenStudy (anonymous):

then plug in the value of 3 into one of the top equations, and you have your x value.

OpenStudy (anonymous):

x=0.5 y=3